# Littlewood-Offord bounds on the symmetric groups and applications

### v1

@@ 9. Further comments
Beyond the applications presented in this note, Theorem 1.20 and Theorem 1.22 also appear to be useful in a variety of counting problems. For instance, Theorem 1.22 is a key tool for establishing strong quantitative invertibility estimates for matrices with fixed row sums and for adjacency matrices of dd–regular digraphs [60, 25]. [^11] More precisely, [60, 25] give sharp lower bounds on the smallest singular value of these random matrices.
To illustrate its application to singularity problems, consider the random 0/11 matrix $\mathcal{Q}_{n,d}$ (introduced in [37]), whose rows are independent vectors containing exactly $d$ ones, where $\min(d,n-d)=\Omega(n)$. We construct $\mathcal{Q}_{n,d}$ row by row. Suppose the first n1n-1 rows are independent and span a hyperplane with normal vector ${\bm{v}}=(v_{1},\ldots,v_{n})$. Conditioned on these rows, the probability that $\mathcal{Q}_{n,d}$ is singular is
@@
where ${\bm{w}}=(w_{1},\ldots,w_{n})$ is the last row. If the LCD of the pair $({\bm{w}},{\bm{v}})$ is large, Theorem 1.22 shows that this probability is small. Tran [60] used this approach to obtain the optimal bound $\exp(-cn)$ for some constant $c>0$. We also note the related work of Jain, Sah, and Sawhney [25] where the authors employed Theorem 1.22 to give a nearly optimal bound for the same problem.
In general, it is expected that if either $\bm{w}$} or $\bm{v}$} arises from a random source, then the LCD is large. We hope to return to this phenomenon in a different venue.
@@ 9.1. Further problems
  • Theorems 1.23 and 1.24 address the problem of determining when $\sup_{x}{\mathbb{P}}(S_{\pi}=x)\geq n^{-3/2}$ or $n^{-5/2}$. Our approach does not, however, yield a seemingly near-optimal inverse result–namely, that if $\sup_{x}{\mathbb{P}}(S_{\pi}=x)$ has order $n^{-1}$, then most of the $w_{i}$ must be zero.
@@
  • While Theorem 1.10 is almost optimal in terms of the size of $Q$, it is interesting to deduce more structure on the sequences $(w_{i})$ and $(v_{i})$ separately. Similarly for Theorems 1.20 and 1.22.
  • While Remark 4.2 shows that Condition 4 is nearly optimal if we rely on Theorem 2.1, our Theorem 1.15 may remain valid without this condition (whereas Theorems 1.17 and 4.4 would require additional assumptions on the $w_{i}$, as stated). It is therefore of interest to remove this condition from Theorem 1.15.
@@
  • It would be interesting to extend Theorems 1.15, 1.17, 4.4, and 1.18 beyond polynomial sequences.
@@
Let ${\mathcal{R}}\subset\{-n,\dots,n\}$ be a set such that $|{\mathcal{R}}|=\Theta_{\delta}(n)$, as defined in Fact 4.1. Let $\Delta$ and $A$ be positive constants, chosen sufficiently large with respect to $\delta$ and $\varepsilon$} (for instance, one may take $\Delta=A$).
@@
$\sup_{x}{\mathbb{P}}\Big(\Big|\sum_{i}n^{3/2-{\varepsilon}}w_{i}\pi(i)-x\Big|\leq\Delta\Big)\ll\int_{|t|\leq 1}\exp\Big\{-\frac{1}{2n^{3}}\sum_{i,j,k,l}\Big\|\frac{tn^{3/2-{\varepsilon}}}{\Delta}(w_{i}-w_{j})(k-l)\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\Big\}\,dt.$
@@
$|{\mathcal{G}}|\geq{\varepsilon}n^{2},\qquad\sum_{(i,j)\in{\mathcal{G}}}(w_{i}-w_{j})^{2}\asymp_{{\varepsilon}}n.$ (52)
@@
$\frac{c_{\delta}}{n^{2}}\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{G}}_{0}\\ r\in{\mathcal{R}}\end{subarray}}\Big\|\frac{tn^{3/2-{\varepsilon}}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{\delta}\frac{1}{n^{2}}\cdot|{\mathcal{G}}_{0}|\cdot n\asymp_{\delta,{\varepsilon}}n.$
@@
$\frac{c_{\delta}}{n^{2}}\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ r\in{\mathcal{R}}\end{subarray}}\Big\|\frac{tn^{3/2-{\varepsilon}}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\gg_{\delta}\frac{1}{n^{2}}\cdot An\log n\cdot n\gg_{\delta}A\log n.$
@@
$\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ r\in{\mathcal{R}}\end{subarray}}\Big\|\frac{tn^{3/2-{\varepsilon}}}{\Delta}(w_{i}-w_{j})r\Big\|^{2}_{{\mathbb{R}}/{\mathbb{Z}}}=\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ r\in{\mathcal{R}}\end{subarray}}\Big|\frac{tn^{3/2-{\varepsilon}}}{\Delta}(w_{i}-w_{j})r\Big|^{2}\asymp_{\delta,\Delta}t^{2}n^{5-2{\varepsilon}}.$
@@
[54, Lemma 3] Let $0<\delta<1$, ${\varepsilon}\leq 10^{-2}\delta$, and let $N$ be an integer with $N\geq 2/\delta$. Suppose that a real number $\theta$ satisfies $\|n\theta\|_{{\mathbb{R}}/{\mathbb{Z}}}\leq{\varepsilon}$ for all $n\in I$, where $I$ is a subset of size at least $\delta N$ of $Q=\{-N,\dots,N\}$. Then there is a natural number $q\leq 2/\delta$ such that
@@
Note that we can give an alternative proof with somewhat implicit constants. Indeed, it is known that if $I\subset Q$, and $|I|\geq\delta|Q|$, then for some sufficiently large $k$ depending on $\delta$, the sumset J=kIkIJ=kI-kI contains a symmetric arithmetic progression $Q^{\prime}=\{-Nd,-(N-1)d,\dots,(N-1)d,Nd\}$ of step $d=O_{\delta}(1)$ and of length $2N+1$. This is an elementary version of the so called Sárkőzy-type theorem in progression, where many more is known (see for instance [52, Lemma 4.4, 5.5] and [53, Lemma B3]). Next, by triangle inequality, for each $n\in J$ we have
$\|n\delta\|\leq 2k{\varepsilon}.$
It thus follows that $\|l(d\theta)\|_{{\mathbb{R}}/Z}\leq 2k{\varepsilon}$, from which we can deduce easily that
$\|d\theta\|_{{\mathbb{R}}/Z}\leq\frac{2k{\varepsilon}}{N}.$
We next deal with polynomials. The following is a version of [54, Proposition 4], where we don’t assume $I$ to be interval.
@@
then there exists a subprogression $Q^{\prime}$ of $Q$ with $|Q^{\prime}|\gg_{d}\delta^{O_{d}(1)}N$ such that $P$ varies by at most $\delta$ on $Q^{\prime}$.
@@
(of Lemma 2.7) To simplify notation we allow implied constants to depend on $d$. We may assume that $\delta\leq c$ for some small constant $c>0$ depending only on $d$. We will also assume that $N$ is sufficiently large.
Applying Proposition B.2, we can find a natural number $q\ll\delta^{-O(1)}$ and an arithmetic subprogression $Q^{\prime}$ of $Q$ such that $|Q^{\prime}|\gg\delta^{O(1)}N$ and such that $P$ varies by at most $\delta$ on $Q^{\prime}$. Writing
@@
for some interval $I^{\prime}\subset Q$ of length $\gg\delta^{O(1)}N$ and some $0\leq r<q$, we conclude that the polynomial $n\mapsto P(qn+r)$ varies by at most $\delta$ on $I^{\prime}$.
Taking $d^{\text{th}}$ order differences, we conclude that the $d^{\text{th}}$ coefficient of this polynomial is $O(\delta^{-O(1)}/N^{d})$; by the binomial theorem, this implies that $n\mapsto P(qn+r)$ differs by at most $O(\delta)$ on $I^{\prime}$ from a polynomial of degree at most d1d-1. Iterating this, we conclude that the $i^{\text{th}}$ coefficient of $n\mapsto P(qn+r)$ is $O(\delta N^{-i})$ for $i=0,\dots,d$, and the claim then follows by inverting the change of variables $n\mapsto qn+r$ (and replacing $q$ with a larger quantity such as $q^{d}$ as necessary). ∎
@@
(of Proposition B.2) We will induct on $d$. The cases $d=1$ is immediate from Lemma B.1. Now suppose that $d\geq 2$, and that the claim had already been proven for d1d-1. It follows from our assumption that
@@
Note that for each $h$, $|\sum_{n\in I_{h}}e(P(n+h)-P(n))|\leq 2N$, we see that for $\gg\delta^{2}N$ choices of $h\in 2Q$ such that
$\frac{1}{N}\Big|\sum_{n\in I_{h}}e(P(n+h)-P(n))\Big|\gg\delta^{2}.$
Note that P(n+h)P(n)P(n+h)-P(n) is a polynomial of degree at most d1d-1 with leading coefficient $h\alpha_{d}n^{d-1}$. We conclude from induction hypothesis (namely Lemma 2.7 for $\alpha_{d}$) that for each such $h$, there exists a natural number $q_{h}\ll\delta^{-O(1)}$ such that
$\|q_{h}h\alpha_{d}\|_{{\mathbb{R}}/{\mathbb{Z}}}\ll\delta^{-O(1)}/N^{d-1}.$
As there are $\gg\delta^{2}N$ choices of $h\in 2Q$ such that the above holds, there are $\gg\delta^{O(1)}N$ integers $n$ in the set $\{-\delta^{-O(1)}N,\delta^{-O(1)}N\}$ such that
$\|n\alpha_{d}\|_{{\mathbb{R}}/{\mathbb{Z}}}\ll\delta^{-O(1)}/N^{d-1}.$
Applying Lemma B.1, we conclude that
$\|q\alpha_{d}\|_{{\mathbb{R}}/{\mathbb{Z}}}\ll\delta^{-O(1)}/N^{d}.$ (54)
Next, we partition $Q$ into arithmetic progressions $Q^{\prime}$ of spacing $q$ and length comparable to $\delta^{C}N$, for a sufficiently large constant $C$ to be chosen. By hypothesis, and by the pigeonhole principle, we have
@@
for at least one such progression $Q^{\prime}$. Assume that $Q^{\prime}=\{i_{0},i_{0}+q,\dots,i_{0}+n^{\prime}q\}$ where $n^{\prime}\asymp\delta^{C}N$. On this progression, for each $0\leq k\leq n^{\prime}$ we write
@@
where $R(k)$ is a polynomial of degree d1d-1 in $k$. Note that by (54), and as $k\leq n^{\prime}\leq\delta^{C}N$ and $kq\leq N$,
$|\alpha_{d}(kq)^{d}|=|(\alpha_{d}q)k(kq)^{d-1}|\ll\delta^{C-O(1)},\forall 0\leq k\leq n^{\prime}.$
to write $\alpha_{d}n^{d}$ as a polynomial in $n$ of degree at most d1d-1, plus an error of size $O(\delta^{C-O(1)})$. We thus can write
@@
for $n\in Q^{\prime}$ for some polynomial $R$ of degree at most d1d-1. Thus, for $C$ large enough, by the triangle inequality we have that
$\frac{1}{|Q^{\prime}|}\Big|\sum_{n\in I\cap Q^{\prime}}e(R(n))\Big|\gg\delta$
and hence by induction hypothesis we may find a subprogression $Q^{\prime\prime}$ of $Q^{\prime}$ of size $|Q^{\prime\prime}|\gg\delta^{O(1)}N$ such that $R$ varies by most $\delta/2$ on $Q^{\prime\prime}$, and thus (for $C$ large enough again) that $P$ varies by at most $\delta$ on $Q^{\prime\prime}$. ∎
@@
$\sup_{x\in{\mathbb{R}}}{\mathbb{P}}\big(|S_{\pi}-x|\leq\delta\big)\ll\int_{|t|\leq 1}\big|\varphi(t/\delta)\big|\,dt\ll\int_{|t|\leq 1}\exp\!\left(-\frac{1}{2n^{3}}\sum_{i,j,k,l}\Big\|\frac{t}{\delta}(w_{i}-w_{j})(v_{k}-v_{l})\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\right)\,dt.$
@@
$\frac{1}{2n^{3}}\,\operatorname{dist}^{2}\!\left(\frac{t}{\delta}{\bm{u}},{\mathbb{Z}}^{n^{4}}\right).$
@@
$\operatorname{dist}\!\left(\frac{t}{\delta}{\bm{u}},{\mathbb{Z}}^{n^{4}}\right)\;\geq\;\min\!\left\{\gamma\left\|\frac{t}{\delta}{\bm{u}}\right\|_{2},\,\kappa\right\}\;\geq\;\min\!\left\{\gamma n^{3/2}\frac{|t|}{\delta},\,\kappa\right\},$
@@
$\sup_{x\in{\mathbb{R}}}{\mathbb{P}}\big(|S_{\pi}-x|\leq\delta\big)\ll\int_{|t|\leq 1}\left(\exp\!\left(-\frac{1}{2}\left(\frac{\gamma t}{\delta}\right)^{2}\right)+\exp\!\left(-\frac{\kappa^{2}}{2n^{3}}\right)\right)\,dt\;\ll\;\frac{\delta}{\gamma}+e^{-\kappa^{2}/2n^{3}}.$

### v2

@@ 12. Further comments
Beyond the applications presented in this note, Theorem 2.11 and Theorem 2.13 also appear to be useful in a variety of counting problems. For instance, Theorem 2.13 is a key tool for establishing strong quantitative invertibility estimates for matrices with fixed row sums and for adjacency matrices of $d$-regular digraphs.
Indeed, in connection with Subsection 3.3, consider the random $0/1$ matrix $Q_{n,d}$ whose rows are independent vectors containing exactly $d$ ones, where $\min(d,n-d)=\Omega(n)$. As in the proof of Theorem 3.9, we construct $Q_{n,d}$ row by row. Suppose the first n1n-1 rows are independent and span a hyperplane with normal vector ${\bm{w}}=(w_{1},\ldots,w_{n})$. Conditioned on these rows, the probability that $Q_{n,d}$ is singular is
@@
where ${\bm{v}}=(v_{1},\ldots,v_{n})$ is the last row. If the LCD of the pair $({\bm{w}},{\bm{v}})$ is large, then Theorem 2.13 implies that this probability is small. Tran [65] used this approach to obtain the optimal bound $\exp(-cn)$ for some constant $c>0$. We also note the related work of Jain, Sah, and Sawhney [28], where the authors employed Theorem 2.13 to give a nearly optimal bound for the same problem.
In general, it is expected that if either $\bm{w}$} or $\bm{v}$} arises from a random source, then the LCD is large. We hope to return to this phenomenon for the matrices introduced in Theorem 3.9 in a different venue.
@@ 12.1. Further problems
  • Theorems 3.2 and 3.3 address the problem of determining when $\sup_{x}{\mathbb{P}}(S_{\pi}=x)\geq n^{-3/2}$ or $n^{-5/2}$. Our approach does not, however, yield a seemingly near-optimal inverse result–namely, that if $\sup_{x}{\mathbb{P}}(S_{\pi}=x)$ has order $n^{-1}$, then most of the $w_{i}$ must be zero.
@@
  • While Theorem 2.1 is almost optimal in terms of the size of $Q$, it is interesting to deduce more structure on the sequences $(w_{i})$ and $(v_{i})$ separately. Similarly for Theorems 2.11 and 2.13.
  • While Remark 6.2 shows that Condition 4 is nearly optimal if we rely on Theorem 4.1, our Theorem 2.6 may remain valid without this condition (whereas Theorems 2.8 and 6.4 would require additional assumptions on the $w_{i}$, as stated). It is therefore of interest to remove this condition from Theorem 2.6.
@@
  • It would be interesting to extend Theorems 2.6, 2.8, 6.4, and 2.9 beyond polynomial sequences.
@@
Let ${\mathcal{R}}\subset\{-n,\dots,n\}$ be a set such that $|{\mathcal{R}}|=\Theta_{\delta}(n)$, as defined in Fact 6.1. Let $\Delta$ and $A$ be positive constants, chosen sufficiently large with respect to $\delta$ and $\varepsilon$} (for instance, one may take $\Delta=A$).
@@
$\sup_{x}{\mathbb{P}}\Big(\Big|\sum_{i}n^{3/2-{\varepsilon}}w_{i}\pi(i)-x\Big|\leq\Delta\Big)=O\Big(\int_{|t|\leq 1}\exp\Big\{-\frac{1}{2n^{3}}\sum_{i,j,k,l}\Big\|\frac{tn^{3/2-{\varepsilon}}}{\Delta}(w_{i}-w_{j})(k-l)\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\Big\}\,dt\Big).$
@@
$|{\mathcal{G}}|\geq{\varepsilon}n^{2},\qquad\sum_{(i,j)\in{\mathcal{G}}}(w_{i}-w_{j})^{2}=\Theta_{{\varepsilon}}(n).$ (54)
@@
$\frac{c_{\delta}}{n^{2}}\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{G}}_{0}\\ r\in{\mathcal{R}}\end{subarray}}\Big\|\frac{tn^{3/2-{\varepsilon}}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta_{\delta}\Big(\frac{|{\mathcal{G}}_{0}|}{n}\Big)=\Theta_{\delta,{\varepsilon}}(n).$
@@
$\frac{c_{\delta}}{n^{2}}\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ r\in{\mathcal{R}}\end{subarray}}\Big\|\frac{tn^{3/2-{\varepsilon}}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Omega_{\delta}(A\log n).$
@@
$\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ r\in{\mathcal{R}}\end{subarray}}\Big\|\frac{tn^{3/2-{\varepsilon}}}{\Delta}(w_{i}-w_{j})r\Big\|^{2}_{{\mathbb{R}}/{\mathbb{Z}}}=\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ r\in{\mathcal{R}}\end{subarray}}\Big|\frac{tn^{3/2-{\varepsilon}}}{\Delta}(w_{i}-w_{j})r\Big|^{2}=\Theta_{\delta,\Delta}(t^{2}n^{5-2{\varepsilon}}).$
@@
[58, Lemma 3] Let $0<\delta<1$, ${\varepsilon}\leq 10^{-2}\delta$, and let $N$ be an integer with $N\geq 2/\delta$. Suppose that a real number $\theta$ satisfies $\|n\theta\|_{{\mathbb{R}}/{\mathbb{Z}}}\leq{\varepsilon}$ for all $n\in I$, where $I$ is a subset of size at least $\delta N$ of $Q=\{-N,\dots,N\}$. Then there is a natural number $q\leq 2/\delta$ such that
@@
Note that one can give an alternative proof with somewhat implicit constants. Indeed, it is known that if $I\subset Q$ and $|I|\geq\delta|Q|$, then for some sufficiently large $k$ depending on $\delta$, the sumset J=kIkIJ=kI-kI contains a symmetric arithmetic progression $Q^{\prime}=\{-Nd,-(N-1)d,\dots,(N-1)d,Nd\}$ with step $d=O_{\delta}(1)$ and length $2N+1$. This is an elementary version of the so-called Sárkőzy-type theorem in progression, for which much more is known (see, for instance, [56, Lemma 4.4, 5.5] and [57, Lemma B3]). Next, by the triangle inequality, for each $n\in J$ we have
$\|n\theta\|_{{\mathbb{R}}/{\mathbb{Z}}}\leq 2k{\varepsilon}.$
It follows in particular that $\|(ld)\theta\|_{{\mathbb{R}}/{\mathbb{Z}}}\leq 2k{\varepsilon}$ for all $|l|\leq N$, from which we easily deduce
$\|d\theta\|_{{\mathbb{R}}/{\mathbb{Z}}}\leq\frac{2k{\varepsilon}}{N}.$
We now turn to polynomials. The following result is a variant of [58, Proposition 4], in which we do not assume $I$ to be an interval.
@@
then there exists a subprogression $Q^{\prime}$ of $Q$ with $|Q^{\prime}|=\Omega_{d}(\delta^{O_{d}(1)}N)$ such that $P$ varies by at most $\delta$ on $Q^{\prime}$.
@@
(of Lemma 4.7) To simplify notation, we allow the implied constants to depend on $d$. We may assume that $\delta\leq c$ some sufficiently small constant $c>0$ depending only on $d$, and that $N$ is sufficiently large.
Applying Proposition B.2, we can find a natural number $q=O(\delta^{-O(1)})$ and an arithmetic subprogression $Q^{\prime}$ of $Q$ such that $|Q^{\prime}|=\Omega(\delta^{O(1)}N)$ and $P$ varies by at most $\delta$ on $Q^{\prime}$. Writing
@@
for some interval $I^{\prime}\subset Q$ of length $|I|=\Omega(\delta^{O(1)}N)$ and some $0\leq r<q$, we conclude that the polynomial $n\mapsto P(qn+r)$ varies by at most $\delta$ on $I^{\prime}$.
Taking $d^{\text{th}}$ order differences, we see that the $d^{\text{th}}$ coefficient of this polynomial is $O(\delta^{-O(1)}/N^{d})$. By the binomial theorem, this implies that $n\mapsto P(qn+r)$ differs by at most $O(\delta)$ on $I^{\prime}$ from a polynomial of degree at most d1d-1. Iterating this argument, we deduce that the $i^{\text{th}}$ coefficient of $n\mapsto P(qn+r)$ is $O(\delta N^{-i})$ for $i=0,\dots,d$. The claim then follows by inverting the change of variables $n\mapsto qn+r$ (and replacing $q$ by a larger quantity such as $q^{d}$ if necessary). ∎
@@
(of Proposition B.2) We argue by induction on $d$. The case $d=1$ follows immediately from Lemma B.1. Now suppose that $d\geq 2$, and that the claim has already been proven for d1d-1. From our assumption,
@@
Since $|\sum_{n\in I_{h}}e(P(n+h)-P(n))|\leq 2N$ for every $h$, there must be $\Omega(\delta^{2}N)$ values of $h\in 2Q$ such that
$\frac{1}{N}\Big|\sum_{n\in I_{h}}e(P(n+h)-P(n))\Big|=\Omega(\delta^{2}).$
Note that P(n+h)P(n)P(n+h)-P(n) is a polynomial of degree at most d1d-1, with leading term $h\alpha_{d}n^{d-1}$. By the induction hypothesis, namely Lemma 4.7, it follows that for each such $h$ there exists a natural number $q_{h}=O(\delta^{-O(1)})$ such that
$\|q_{h}h\alpha_{d}\|_{{\mathbb{R}}/{\mathbb{Z}}}=O(\delta^{-O(1)}/N^{d-1}).$
Since there are $\Omega(\delta^{2}N)$ choices of $h\in 2Q$ for which this holds, we obtain $\Omega(\delta^{O(1)}N)$ integers $n$ in the interval $\delta^{-O(1)}N,\delta^{-O(1)}N]$ such that
$\|n\alpha_{d}\|_{{\mathbb{R}}/{\mathbb{Z}}}=O(\delta^{-O(1)}/N^{d-1}).$
Applying Lemma B.1, we conclude
$\|q\alpha_{d}\|_{{\mathbb{R}}/{\mathbb{Z}}}=O(\delta^{-O(1)}/N^{d}).$ (56)
Next, we partition $Q$ into arithmetic progressions $Q^{\prime}$ of spacing $q$ and length comparable to $\delta^{C}N$, for a sufficiently large constant $C$ to be chosen. By the hypothesis and an application of the pigeonhole principle, we have
@@
for at least one such progression $Q^{\prime}$. Assume that $Q^{\prime}=\{i_{0},i_{0}+q,\dots,i_{0}+n^{\prime}q\}$ where $n^{\prime}=\Theta(\delta^{C}N)$. On this progression, for each $0\leq k\leq n^{\prime}$ we write
@@
where $R(k)$ is a polynomial of degree d1d-1 in $k$. Since $k\leq n^{\prime}=\Theta(\delta^{C}N)$ and $kq\leq N$, (56) implies
$|\alpha_{d}(kq)^{d}|=|k(kq)^{d-1}(q\alpha_{d})|=O(\delta^{C-O(1)}).$
Thus for $n\in Q^{\prime}$ we may write
@@
for some polynomial $R$ of degree at most d1d-1. Choosing $C$ sufficiently large and applying the triangle inequality, we obtain
$\frac{1}{|Q^{\prime}|}\Big|\sum_{n\in I\cap Q^{\prime}}e(R(n))\Big|=\Omega(\delta).$
By the induction hypothesis, there exists a subprogression $Q^{\prime\prime}$ of $Q^{\prime}$ of size $|Q^{\prime\prime}|=\Omega(\delta^{O(1)}N)$ on which $R$ varies by at most $\delta/2$. Taking $C$ sufficiently large, it follows that $P$ varies by at most $\delta$ on $Q^{\prime\prime}$. ∎
@@
$\sup_{x\in{\mathbb{R}}}{\mathbb{P}}\big(|S_{\pi}-x|\leq\delta\big)=O\Big(\int_{|t|\leq 1}\big|\varphi(t/\delta)\big|\,dt\Big)=O\Big(\int_{|t|\leq 1}\exp\Big(-\frac{1}{2n^{3}}\sum_{i,j,k,l}\Big\|\frac{t}{\delta}(w_{i}-w_{j})(v_{k}-v_{l})\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\Big)\,dt\Big).$
@@
$\frac{1}{2n^{3}}\,\operatorname{dist}^{2}\Big(\frac{t}{\delta}{\bm{u}},{\mathbb{Z}}^{n^{4}}\Big).$
@@
$\operatorname{dist}\Big(\frac{t}{\delta}{\bm{u}},{\mathbb{Z}}^{n^{4}}\Big)\geq\min\Big\{\gamma\Big\|\frac{t}{\delta}{\bm{u}}\Big\|_{2},\,\kappa\Big\}\geq\min\Big\{\gamma n^{3/2}\frac{|t|}{\delta},\,\kappa\Big\},$
@@
$\sup_{x\in{\mathbb{R}}}{\mathbb{P}}\big(|S_{\pi}-x|\leq\delta\big)=O\Big(\int_{|t|\leq 1}\Big(\exp\Big(-\frac{1}{2}\Big(\frac{\gamma t}{\delta}\Big)^{2}\Big)+\exp\Big(-\frac{\kappa^{2}}{2n^{3}}\Big)\Big)\,dt\Big)=O\Big(\frac{\delta}{\gamma}+e^{-\kappa^{2}/2n^{3}}\Big).$